OpenStudy (anonymous):

Find the mean and standard deviation of each data set. Round to the nearst hundredth. 1. 4 8 5 12 3 9 5 2

OpenStudy (anonymous):

i already know the mean i just dont know how to find the standard deviation.

jimthompson5910 (jim_thompson5910):

You should have found that the mean is 6. Now subtract the sample mean from EVERY data value: 4-6 = -2 8-6 = 2 5-6 = -1 12-6 = 6 3-6 = -3 9-6 = 3 5-6 = -1 2-6 = -4

jimthompson5910 (jim_thompson5910):

So the differences are: -2, 2, -1, 6, -3, 3, -1, -4

jimthompson5910 (jim_thompson5910):

Now square each difference: (-2)^2 = 4 (2)^2 = 4 (-1)^2 = 1 (6)^2 = 36 (-3)^2 = 9 (3)^2 = 9 (-1)^2 = 1 (-4)^2 = 16

jimthompson5910 (jim_thompson5910):

So we have the squares 4, 4, 1, 36, 9, 9, 1, 16

jimthompson5910 (jim_thompson5910):

Now add up each square: 4+4+1+36+9+9+1+16 = 80

jimthompson5910 (jim_thompson5910):

Next, divide that sum by n-1 = 8-1 = 7 to get 80/7 = 11.4285714285714 So the sample variance is roughly 11.4285714285714 To find the sample standard deviation, just take the square root of the sample variance.

OpenStudy (anonymous):

Oooh... Thanks! I think I understand now

jimthompson5910 (jim_thompson5910):

Yes, there are a lot of steps, but that's great.

jimthompson5910 (jim_thompson5910):

To speed things up, you can just use a calculator (which should be allowed)

OpenStudy (anonymous):

KK :)